OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x? (−2, 0) and (3, 0) (−2, 0) and (−3, 0) (2, 0) and (−3, 0) (2, 0) and (3, 0) I have an idea of how to solve this, I think it would be the x = -b/2(a) formula then to plug in, but I am at a complete loss.

OpenStudy (zpupster):

no subtract 6 both sides y = x2 - x - 6 factor the polynomial and then set y to zero

OpenStudy (anonymous):

So i'm subtracting six from itself and x?

OpenStudy (zpupster):

like this y + 6 -6= x2 - x -6

OpenStudy (anonymous):

For finding \(x\)-intercept, always put \(y=0\), the value of \(x\) you get is your \(x\)-intercept.. :)

OpenStudy (anonymous):

Set \(y=0\) and find \(x\)..

OpenStudy (anonymous):

Can you show me how to do that?

OpenStudy (anonymous):

Sure : \(y + 6 = x^2 - x\) Set \(y = 0\) : \(0 + 6 = x^2 - x\)

OpenStudy (anonymous):

\[x^2 - x = 6 \implies \color{green}{x^2 - x - 6 = 0}\]

OpenStudy (anonymous):

Can you solve this quadratic equation and find \(x\)??

OpenStudy (anonymous):

Not really, I am writing every thing you say down, but I am stumped

OpenStudy (anonymous):

Haven't read Quadratic Formula?

OpenStudy (anonymous):

Okay, tell me where are you having doubt?

OpenStudy (anonymous):

Just, how to set this one up, I just feel so overwhelmed by it for some reason

OpenStudy (anonymous):